OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (phi):

you need to find u/|u|

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

is the absolute value of u <4, 3>?

OpenStudy (phi):

find |u| |u| = \( \sqrt{ u \cdot u} \) so first find u dot u can you do that ?

OpenStudy (anonymous):

how exactly would I do that? just multiply u by u?

OpenStudy (phi):

the "dot product" of a vector means 1) multiply "corresponding" elements 2) add up the products example: < a,b> dot < c,d> = a*c + b*d

OpenStudy (anonymous):

so what would be the c,d part? I only have a,b

OpenStudy (phi):

you are doing c dot c

OpenStudy (anonymous):

okay, so it'd be -4*-4+-3*-3?

OpenStudy (phi):

yes

OpenStudy (anonymous):

16+9=25 so sqrt(u dot u) is 5?

OpenStudy (phi):

yes. that means the "length" of u is 5 or |u| = 5 now divide u by 5 (each element of 5) to make it a "unit vector"

OpenStudy (phi):

** now divide u by 5 (each element of u by 5) to make it a "unit vector"

OpenStudy (anonymous):

\[<\frac{ -4 }{ 5 },\frac{ -3 }{ 5 }>\]

OpenStudy (anonymous):

@phi is that right?

OpenStudy (phi):

yes, or < -0.8, -0.6> notice if we find its length we get \[ \sqrt{ -0.8\cdot -0.8 + -0.6\cdot -0.6 }= \sqrt{0.64+.36}= \sqrt{1}= 1 \] the new vector has unit length (i.e. its length is one)